OpenStudy (anonymous):

prove that; sin3(theta)=3sin(theta)-4sin^3(theta)

OpenStudy (anonymous):

ill write x for theta sin3x = sin(2x + x) = sin2x cos x + cos2xsinx = 2sinxcos^2x + sinx (1- 2sin^2x) = sinx (2cos^2x + 1 - 2sin^2x) = sinx (2(1 - sin^2x) +1-2sin^2x) = 3sinx - 4sin^3x sin 3(theta) = 3sin(theta) - 4sin^3(theta)

OpenStudy (anonymous):

sin 3s